OpenStudy (anonymous):

Simplify (-2^4) x^3 y^6 _______________ (-2^2) x^-2 y^-6

OpenStudy (anonymous):

\[\frac{16x^3y^6}{\frac{4}{x^2y^6}}\] \[16x^3y^6*\frac{x^2y^6}{4}\] \[4x^5y^12\]